OpenStudy (anonymous):

eight years ago, a man was four years less than three times as old as his son's age. six years from now, he will be twice as old as his son. how old are they now?

OpenStudy (anonymous):

Let in present age of the son is x, and age of father is y now 8 year ago age of son and father is x-8 and y-8 so relation bw father and son ages 8 year ago is y-8=3(x-8)-4 now after 6 year from now equation will be y+6=2(x+6) solve these two eqn you will get the value of x and y.

OpenStudy (anonymous):

Mans age = x sons age = y 8 years ago 3(y-8) -4 = x-8 six years from now 2(y+6) = x+6 solving y=30 x= 70

OpenStudy (anonymous):

in certain office, betty can type twice as many words per minute as annie. cindy can type 80 more words tha annie. if the three can type 600 words per minute. how fast can each girl type?

OpenStudy (anonymous):

betty=x annie=y cindy=z x=2y z=80 + y x+y+z=600 solving y=130 x= 260 z= 210

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

jake is nine years older than leth. three years ago, jake was twice as old as leth, how old are the two?